OpenStudy (anonymous):

1 hemoglobin molecule contains 4 Ferum atoms.If 1g of hemoglobin contains 0.0341g of Ferum, calculate the relative molecular mass of hemoglobin.... PLS HELP!!!!!

OpenStudy (aaronq):

make a ratio with what you know: \(\dfrac{1~g~Hb}{0.0341~g~Fe}=\dfrac{M_{Hb}}{4*M_{Fe}}\) the molar mass of Fe is 55.845 g/mol \(\dfrac{1~g~Hb}{0.0341~g~Fe}=\dfrac{M_{Hb}}{4*55.845 ~g/mol}\) \(M_{Hb}=\dfrac{1~g~Hb*[4*55.845 ~g/mol]}{0.0341~g~Fe}=6550.73~g/mol\)

OpenStudy (aaronq):

Although, i think the mass of Fe the question should be 0.00341 g, because otherwise the final answer (above) is incorrect by an order of magnitude.